The graphs of the equations 4x-y=6 and x+y=4 intersect at the point whose coordinates are
1.(2,-2)
2.(5,-1)
3.(1,3)
4.(2,2)

Answer by nerdybill(7384)

Solve the system of equations.
Do this by the addition method -- add both equations together:
4x-y=6
x+y=4
-------
5x =10
x = 2
.
Substitute the above into:
x+y=4
2+y=4
y=2
.
Therefore, the graphs cross at:
(x,y) = (2,2)